A hanging spring is compressed 3 inches from its rest position and released at t = 0 seconds. It returns to the same position after 0.8 seconds.
Find:
a) the amplitude of the motion

b) the period of the motion

c) the frequency of the motion

d) a function that models the displacement, y, of the end of the spring from the rest position at time, t.

e) the displacement from the rest position at t= 3 min rounded to the tenths place

Answer by Alan3354(69443)

A hanging spring is compressed 3 inches from its rest position and released at t = 0 seconds. It returns to the same position after 0.8 seconds.
Find:
a) the amplitude of the motion
3 inches
-----
b) the period of the motion
1.6 seconds
----
c) the frequency of the motion
1/1.6 = 5/8 Hz
---
d) a function that models the displacement, y, of the end of the spring from the rest position at time, t.
y = 3cos(16t/(5pi))
--------
e) the displacement from the rest position at t= 3 min rounded to the tenths place
y = 3cos(16t/(5pi))
y @ 180 seconds = y = 3cos(16*180/(5pi)) = 3cos(576/pi) = 1.26879
--> 1.3 inches